# Chapter 0 Hadron Production Processes

### v1

@@ 2 Charm and Beyond
As early as in 1964, the idea came up that there might be more flavors than $\{u,d,s\}$. James Bjorken (1934-2024) and Sheldon Glashow (born 1932) hypothesized the existence of another quark, which they named charm ($c$). As Glashow declared in 1976 in a newspaper article in New York Times the charm quark received its name because of the symmetry it brought to the Subnuclear world as cited in [35]. Experimentally, a signature of charm was observed first in 1974 in two independent experiments at BNL by Samuel C. C. Ting (born 1936) and collaborators [36] and at SLAC by the group of Burton Richter (1931-2018) [37]. Both teams found at an energy of $3.1~{}\mathrm{GeV}$ a narrow resonance with a width of only $5~{}\mathrm{MeV}$. A particular strong confirmation for the existence of charm was that the BNL group used the purely hadronic nuclear reaction $pBe\to e^{+}e^{-}+X$ while the SLAC group observed the new resonance in purely leptonic $e^{+}e^{-}$ annihilation reactions. Thus proving the ubiquity of the phenomenon. The BNL group gave that state the name $J$, the competitors from SLAC chose $\psi$, and finally, as a compromise and honoring the work of both teams, the two labels emerged into the official name $J/\psi$.
@@
Ting and Richter met on November 15, 1974; they issued a press release about their discoveries. The charm story then took up speed: On November 21, 1974, the SPEAR@SLAC collaboration already reported the observation of $\psi^{\prime}(3700)$, an excited state of the $J/\psi(3100)$. Both particles are charmonium states, composed of $c\bar{c}$ valence quarks. The modern spectroscopic notation is $J/\psi(1S)$ and $\psi^{\prime}(2S)$, following the constituent quark model and indicating that $\psi^{\prime}$ is understood as a radial (compressional) excitation of $J/\psi$. The latest mass values are $M_{J/\psi}=3096.900\pm 0.006~{}\mathrm{MeV}$ and $M_{\psi^{\prime}}=3686.097\pm 0.010~{}\mathrm{MeV}$ [38].
@@
The two group leaders Ting and Richter were awarded the 1976 Nobel prize, honoring the great achievement of the two teams for for discovering the subatomic $J/\psi$ particle. Their epochal discoveries became known as the November Revolution of particle physics.
@@
In 1986, Matsui and Satz predicted $\ldots$ that $J/\psi$ suppression in nuclear collisions should provide an unambiguous signature of quark-gluon plasma formation [39]. Since then, the $J/\psi$ plays a special role in the QGP searches in heavy-ion collisions at ultra-relativistic energies at RHIC and the LHC.
@@
A multitude of hidden charm charmonium states and open charm mesons and baryons have been observed, containing a charmed quark or anti-quark to which $\{u,d,s\}$ partners and the respective anti-quarks are attached. There is strong evidence that charmed hadrons may develop exotic configurations, e.g., tetra-quark states or molecular-like states formed by D mesons. These so-called $X,Y,Z$ states fall out of the systematics of the otherwise very successful quark models for quarkonia. Their spectroscopy is discussed elsewhere in this Encyclopedia [40].
@@
Corresponding hadron configurations are found for bottom ($b$) quarks and have to be expected in the top ($t$) quark sector. Bottom and top quarks together represent the third generation of the quark family. In Fig. 5, a table summarizing the three quark generations and their properties is shown.
@@ 1 Aspects of QCD
QCD is the fundamental theory of strong interactions and thus of hadrons666 Lev Borisovich Okun (1929-2015) introduced the term hadron (ancient Greek for heavy or strong) in a plenary talk at the 1962 International Conference on High Energy Physics.. The beauty and peculiarities of QCD as a non-Abelian quantum gauge field theory of intrinsically non-perturbative character will be discussed elsewhere in this Encyclopedia. Here, an overview is given on the aspects that are of special interest and relevance for hadron physics and the production of hadrons.
@@
In the early days of nuclear and elementary particle physics, the plethora of particles observed in cosmic-ray and high energy experiments made improbable their being elementary particles. Obviously, a new approach was necessary, allowing for bringing order into the data. In 1961, Yuval Ne’eman (1925-2006) introduced the classification of hadrons through the SU(3) flavor symmetry, exploiting the representation of the SU(3) group in terms of eight linearly independent 3-by-3 matrices and three elementary degrees of freedom, later named the Eightfold Way [41, 42]. Similar group-theoretical descriptions were invented independently by Murray Gell-Mann (1929-2019), who finally formulated QCD [43]. Murray Gell-Mann was awarded the Nobel Prize in Physics in 1969 for his contributions and discoveries concerning the classification of elementary particles and their interactions.
@@
In parallel to and independently of Ne’eman and Gell-Mann, George Zweig (born 1937) developed during his stay as a post-doc at CERN his own version of a SU(3) scheme of strongly interacting particles [44, 45]. His contributions to QCD are honored in the OZI rule, accounting, e.g., for the hindered decay of the $\varphi(1020)$ vector meson. Thus, by 1964, two competing group theoretical approaches were available, relating the abstract mathematical structure of non-commuting Lie algebra (named after the mathematician Sophus Lie (1842-1899)) to elementary particle physics. Although at that time the authors insisted on essential differences between their theories, there were obvious overlapping similarities. Both approaches were in need of introducing three elementary entities in order to place mesons and baryons on the same footing. While Zweig chose the name aces, inspired by the joker cards of card games, Gell-Mann opted for quarks [^7] Three quarks for Muster Mark” is a line from James Joyce’s novel “Finnegans Wake.” (It took for James Augustine Aloysius Joyce (1982-1941) 16 years to complete this experimental novel.) The phrase is notable because it provided the name “quark” for a fundamental subatomic particle, chosen by Murray Gell-Mann. Gell-Mann initially considered the word “kwork” before encountering the Joyce passage and finding it more fitting. While the original line seems to rhyme “quark” with “Mark,” Gell-Mann proposed a justification for pronouncing it ”kwork” by associating it with the pub owner’s call of “Three quarts for Mister Mark. - which finally was adopted by the community. In Figure (Table) 5, the quarks and their properties are listed. The discovery of new particle species ($J/\psi$) in 1974 enforced the introduction of the charm (c) quark and later discoveries of particles required to extend the list of QCD flavors to bottom (b) and top (t) quarks.
Figure 5: Properties of quarks of the first $(u,d)$, second $(c,s)$, and third $(t,b)$ generation [38]. As a generalization of isospin, originally introduced for $\{u,d\}$ systems and used above to distinguish proton and neutrons by the eigenvalues of the third component of the isospin operator $I_{3}=\pm\frac{1}{2}$, heavier quarks are characterized by the flavors $c,s,t,b$ with values $0,1$. In this chapter, only production processes involving hadrons built of the first and second quark generation are considered. The figure is adapted from Ref. [46].
The works of Yuval Ne’eman, Murray Gell-Mann, and George Zweig were the long-awaited breakthrough towards a new understanding of hadrons in terms of a few elementary degrees of freedom, given by the massive quark Fermion fields and by gluon gauge vector fields, which are the force carriers of strong interactions. Different from the photons of QED, gluons carry color charges and interact with each other.
An immediate consequence of the group-theoretical approach is to introduce order schemes in terms of multiplets which are defined by the invariants of the respective mathematical group. In Fig. 6, the lowest octet and decuplet SU(3) multiplets of baryons, composed of up (u), down (d), and strange (s) quarks, are shown by being arranged in a diagram constructed by the two group invariants hypercharge, $Y=B+S$, defined by the baryon number $B$ and strangeness $S$ and the third component of the SU(2) isospin operator, $I_{3}=Q-\frac{1}{2}Y$, where $Q$ is the baryon charge in units of the elementary charge $e$. The algebra underlying multiplet diagrams as in Fig. 6 was described by de Swart [47] (Johan de Swart (1931-2014)) in a comprehensible, compact manner and has been used widely until today.
@@
Although group theory does not provide by itself absolute mass scales, the mathematical structures predict highly valuable sum rules relating the masses within multiplets. In [48], Murray Gell-Mann exploited those relations to identify the already known two-pion rho-meson ($\rho$) resonance to be part of a vector meson octet. He also predicted that there must be an isoscalar partner of the rho-meson decaying into three pions, which he named omega-meson ($\omega$). Moreover, at the International Conference on High-Energy Nuclear Physics, Geneva, 1962, Gell-Mann used the mass relations to predict that the still missing last decuplet member of strangeness $S=-3$, i.e., the $\Omega^{-}$ hyperon, would have the mass $M^{(theo)}_{\Omega}=1676~{}\mathrm{MeV}$. Two years later, the particle was indeed found at Brookhaven by Barnes et al. [49, 50] at a mass of $M^{(exp)}_{\Omega}=1674\pm 3~{}\mathrm{MeV}$ [^8] The BNL 80-in. hydrogen bubble chamber was exposed to a mass-separated beam of $5.0~{}\mathrm{BeV/c}$ K mesons at the Brookhaven AGS. About 100,000 pictures were taken containing a total K track. The analysis shown a single event associated with the $\Omega^{-}\to nK^{0}\pi^{-}$ decay.. That extraordinarily good agreement gave strong support and confidence to the group-theoretical SU(3) approach.
@@
At the same time as Gell-Mann but independently, Susumo Okubo (1930-2015) was pursuing at Rochester his own studies of unitary symmetry in strong interactions, leading to mass relations of the same type [41, 51, 52]. Since then, the SU(3) mass relations are known as Gell-Mann-Okubo (GMO) mass formulas [^9] Deviations from the GMO mass formula predictions are often important hints for mass shifts form dynamical mixing..
@@
One of the central predictions of QCD is the parton structure of hadrons. Once that conjecture was confirmed by experiment in the early 1970’s [53], QCD evolved into the nowadays accepted standard model of strong interaction physics. QCD describes hadrons as systems of quarks, electrically charged spin-$\frac{1}{2}$ fermions with flavor, color and isospin, and gluons, colored spin-$1$, self-interacting vector gauge fields. Outstanding features of QCD are quark confinement and color-neutrality of physical particles, running coupling constants, and asymptotic freedom. Renormalization group techniques and regularization methods are indispensable parts of the QCD mathematics toolbox. Spontaneous symmetry breaking is a defining mechanism of QCD. The complexities of the non-Abelian structures of QCD (named after Niels Henrik Abel (1802-1829)), leading to higher order non-linear self-interactions, inhibit solving the deeply nested field equations in closed form. Only in the ultra-high energy regime of asymptotic freedom, where perturbative treatments become possible, leading order results can be derived analytically. The electro-weak theory of Salam, Weinberg, and Glashow, and QCD are the constituencies of the Standard Model (SM) of elementary particle physics. Mohammad Abdus Salam (1926-1996) shared the 1979 Nobel Prize in Physics with Steven Weinberg (1933-2021) and Sheldon Glashow (born 1932) for his contribution to the electroweak unification theory.
For a long time, the missing link was a mechanism which could explain within the SM the origin of mass. In 1964, Peter Higgs (1929-2024) proposed a mechanism for mass generation which was compatible with the SM but postulated the existence of a boson (named after Satyendra Nath Bose (1894-1974)) multiplet [54, 55]. In 2012, finally the (neutral) Higgs boson $H^{0}$ (colloquially called “God Particle”) was observed at about the expected mass $m_{H}=(125.20\pm 0.11)~{}\mathrm{GeV}$ [38] at the LHC at CERN. Since then, the SM is a closed theory, although lacking a link to gravitation. For this work, Higgs received the Nobel Prize in Physics in 2013, which he shared with François Englert (born 1932).
@@
Figure 6: The first two baryon SU(3) flavor multiplets are given by an octet (left) and a decuplet (right). The valence quark content of the baryons is indicated explicitly. The vertical axes are representing the hypercharge $Y=S+B$, given by the strangeness $S$ and baryon number $B$; the horizontal axes indicate the third component $I_{3}=Q-\frac{1}{2}Y$ of the isospin $I$ includes also the charge number $Q$. The group theoretical background and construction of these kind of diagrams is discussed in depth in textbooks, see, e.g.[56].
@@
Mesons and baryons may be considered as the asymptotic mass eigenstates of QCD in the low energy limit. However, they are of fundamentally different character. Considered as systems of quarks and gluons, mesons are formed by an even number of valence quarks, while baryons are defined as the particles containing an odd number of valence quarks. In both cases, an arbitrary amount of gluons is allowed in addition. Since quarks are Fermions of spin $s=\frac{1}{2}$ and positive parity, baryons carry half-integer spin quantum numbers $J_{B}=\frac{1}{2},\frac{3}{2},\frac{5}{2},\ldots$, while mesons have integer spin quantum numbers $J_{M}=0,1,2\ldots$. Moreover, baryons play a special role in cosmology because they account as protons and neutrons together with a minor contribution of electrons for the matter content of the Universe.
@@
Figure 7: The lowest meson SU(3) flavor nonets of pseudo-scalar (left), vector (center), and scalar mesons (right) as functions of hypercharge Y and isospin I3. Different to the baryons, mesonic two-quark systems of $\{u,d,s\}\otimes\{\bar{u},\bar{d},\bar{s}\}$ structure arrange into SU(3) octets and accompanying singlets. The isoscalar (charge neutral) octet mesons obtain an isoscalar singlet partner. The singlet mesons are included into the octet diagrams, giving rise to the octet-singlet pairs $\{\eta,\eta^{\prime}\}$, $\{\omega,\phi\}$, and $\{\sigma\simeq f_{0}(500),\sigma^{\prime}\simeq f_{0}(980)\}$ in the pseudo-scalar, vector, and scalar meson sectors, respectively. Singlet and octets are mixing by, in principle, arbitrary mixing angles. A widely used choice is Ideal Mixing, chosen such that the isoscalar octet mesons $\eta,\omega,\sigma$ do not contain a $s\bar{s}$ component. Electromagnetic effects lead to $\omega-\rho$ mixing and correspondingly $\sigma-\delta^{0}$ mixing has to be expected. Scalar mesons may contain multi-pion. i.e., multi-$q\bar{q}$ components. The complexities of scalar mesons are especially visible in the isovector scalar meson $\delta\simeq a_{0}(980)$. $a_{0}(980)$ is decaying preferentially into $\eta\pi$, finally ending in three pions, but also $K\bar{K}$ and $\eta^{\prime}\pi$ decays are reported [38]. The physical counterpart of the $\kappa$ meson is the $K^{\ast}_{0}(700)$ meson of a notoriously uncertain pole structure [38]. Thus, the physically observed mesons are not identical with the bare SU(3) mesons.

### v2

@@ 2 Charm and Beyond
As early as 1964, the idea arose that there might be more flavors than $\{u,d,s\}$. James Bjorken (1934-2024) and Sheldon Glashow (born 1932) hypothesized the existence of another quark, which they named charm ($c$). As Glashow declared in 1976 in a newspaper article in New York Times the charm quark received its name because of the symmetry it brought to the Subnuclear world, as cited in [38]. Experimentally, a signature of charm was first observed in 1974 in two independent experiments at BNL by Samuel C.C. Ting (born 1936) and collaborators [39] and at SLAC by the group of Burton Richter (1931-2018) [40]. Both teams found a narrow resonance at an energy of $3.1~\mathrm{GeV}$ with a width of only $5~\mathrm{MeV}$. A particularly strong confirmation of the existence of charm was that the BNL group used the purely hadronic nuclear reaction $p\mathrm{Be}\to e^{+}e^{-}X$, while the SLAC group observed the new resonance in purely leptonic $e^{+}e^{-}$ annihilation reactions. Thus, this proved the ubiquity of the phenomenon. The BNL group gave that state the name $J$, the competitors from SLAC chose $\psi$, and finally, as a compromise and in honor of the work of both teams, the two labels merged into the official name $J/\psi$.
@@
Ting and Richter met on November 15, 1974; they issued a press release about their discoveries. The charm story then took off: On November 21, 1974, the SPEAR@SLAC collaboration already reported the observation of $\psi^{\prime}(3700)$, an excited state of the $J/\psi(3100)$. Both particles are charmonium states, composed of $c\bar{c}$ valence quarks. The modern spectroscopic notation is $J/\psi(1S)$ and $\psi^{\prime}(2S)$, following the constituent quark model and indicating that $\psi^{\prime}$ is understood as a radial (compressional) excitation of $J/\psi$. The latest mass values are $M_{J/\psi}=3096.900\pm 0.006~\mathrm{MeV}$ and $M_{\psi^{\prime}}=3686.097\pm 0.010~\mathrm{MeV}$ [41].
@@
The two group leaders, Ting and Richter, were awarded the 1976 Nobel Prize, honoring the great achievement of the two teams for discovering the subatomic $J/\psi$ particle. Their epochal discoveries became known as the November Revolution in particle physics.
@@
In 1986, Matsui and Satz predicted $\ldots$ that $J/\psi$ suppression in nuclear collisions should provide an unambiguous signature of quark-gluon plasma formation [42]. Since then, $J/\psi$ has played a special role in searching for QGP in heavy-ion collisions at ultra-relativistic energies at RHIC and the LHC.
@@
A multitude of hidden charm charmonium states and open charm mesons and baryons have been observed, containing a charmed quark or anti-quark to which $\{u,d,s\}$ partners and the respective anti-quarks are attached. There is strong evidence that charmed hadrons may develop exotic configurations, e.g., tetra-quark states or molecular-like states formed by D-mesons. These so-called $X,Y,Z$ states fall out of the systematics of the otherwise very successful quark models for quarkonia. Their spectroscopy is discussed elsewhere in this Encyclopedia [43].
As an explanation for the still unsatisfactorily solved CP violation puzzle, Makoto Kobayashi and Toshihide Maskawa predicted in 1973 the existence of a third family of quarks [33]. That conjecture was partly confirmed in 1977 by the discovery of the bottom ($b$) or beauty quark [44] and completed by the observation of the top ($t$) quark in 1995 [45, 46]. Both the bottom and top quarks were first detected at FNAL.
The Nobel Prize of 2008 was awarded one half to Yoichiro Nambu (1921-2015) for the discovery of the mechanism of spontaneous broken symmetry in subatomic physics, the other half was shared by Makoto Kobayashi (born 1944) and Toshihide Maskawa (1940-2021) for the discovery of the origin of the broken symmetry which predicts the existence of at least three families of quarks in nature. Bottom and top quarks together represent the third generation of the quark family.
@@
Charmonium-like hadron configurations have been found for bottom quarks ($\Upsilon$ family) and were also expected to exist in the top ($t$) quark sector. The latter is indeed confirmed by the first observation of a $t\bar{t}$ vector meson by the CMS experiment at the LHC [47]. In Fig. 5, a table summarizing the three quark generations and their properties is shown.
@@ 1 Aspects of QCD
QCD is the fundamental theory of strong interactions and thus of hadrons [^6] Lev Borisovich Okun (1929-2015) introduced the term hadron (ancient Greek for heavy or strong) in a plenary talk at the 1962 International Conference on High Energy Physics.. The beauty and peculiarities of Quantum Chromodynamics (QCD) as a non-Abelian quantum gauge field theory of intrinsically non-perturbative character will be discussed elsewhere in this Encyclopedia. Here, an overview is provided of the aspects that are of special interest and relevance for hadron physics and the production of hadrons.
@@
In the early days of nuclear and elementary particle physics, the plethora of particles observed in cosmic-ray and high-energy experiments made it improbable that they were elementary particles. Obviously, a new approach was necessary, allowing one to bring order to the data. In 1961, Yuval Ne’eman (1925-2006) introduced the classification of hadrons by means of the SU(3) flavor symmetry group, exploiting the representation of the SU(3) group in terms of eight linearly independent 3-by-3 matrices and three elementary degrees of freedom. At the same time, Murray Gell-Mann (1929-2019) also independently pursued similar group-theoretical research and named the approach the Eightfold Way in a book from 1964 [48, 49].
@@
In parallel to Ne’eman and Gell-Mann, George Zweig (born 1937) developed his own version of a SU(3) scheme of strongly interacting particles [50, 51], mostly during his stay as a post-doctoral researcher at CERN. His contributions to QCD are engraved in the OZI rule, accounting, e.g., for the hindered decay of the $\varphi(1020)$ vector meson. Thus, by 1964, two competing group theoretical approaches were available, relating to the abstract mathematical structure of a non-commuting Lie algebra (named after the mathematician Sophus Lie (1842-1899) to elementary particle physics. Although at that time, the authors insisted on essential differences between their theories, there were obvious overlapping similarities. Both approaches needed three elementary entities in order to place mesons and baryons on the same footing. Zweig chose the name aces, inspired by the joker cards of card games. However, in a paper of 1964 [52], Gell-Mann introduced A Schematic Model of Baryons and Mesons. That paper can be considered foundational to the theory of elementary particle physics: Gell-Mann points out that A simpler and more elegant scheme can be constructed if we allow non-integral values for the charges. He also denotes the constituents of his model as quarks by explicitly referring to the novel Finnegans Wake by James Joyce, and that naming was finally adopted by the community.  [^7] Three quarks for Muster Mark” is a line from James Joyce’s novel “Finnegans Wake.” (It took James Augustine Aloysius Joyce (1982-1941) 16 years to complete this experimental novel.) The phrase is notable because it provided the name “quark” for a fundamental subatomic particle, chosen by Murray Gell-Mann. Gell-Mann initially considered the word “kwork” before encountering the Joyce passage and finding it more fitting. While the original line seems to rhyme “quark” with “Mark,” Gell-Mann proposed a justification for pronouncing it “kwork” by associating it with the pub owner’s call of “Three quarts for Mister Mark. In Figure (Table) 5, the quarks and their properties are listed. Murray Gell-Mann received the Nobel Prize in Physics in 1969 for his contributions and discoveries concerning the classification of elementary particles and their interactions.
Group theory predicts highly valuable ordering schemes by which particles of equal symmetry class can be arranged in multiplets. Graphical representations of the lowest baryon octet and decuplet are shown in Fig.6. The pseudo-scalar, scalar, and vector meson nonets are displayed in Fig. 7.
@@
Although group theory does provide neither absolute mass scales nor absolute (numerical) values of coupling constants, relations among coupling constants and sum rules for the masses within a given multiplet can be derived, typically reducing the problem to the determination of a few free parameters. In [53], Murray Gell-Mann exploited those relations to identify the already known two-pion rho-meson ($\rho$) resonance as part of a vector meson octet. He also predicted that there must be an isoscalar partner of the rho-meson decaying into three pions, which he named the omega-meson ($\omega$). Moreover, at the International Conference on High-Energy Nuclear Physics, Geneva, 1962, Gell-Mann used the mass relations to predict that the still missing last decuplet member of strangeness S=3S=-3, i.e., the $\Omega^{-}$ hyperon, would have the mass $M^{(theo)}_{\Omega}=1676~\mathrm{MeV}$. Two years later, the particle was indeed found at Brookhaven by Barnes et al. [54, 55] at a mass of $M^{(exp)}_{\Omega}=1674\pm 3~\mathrm{MeV}$ [^8] The BNL 80-in. hydrogen bubble chamber was exposed to a mass-separated beam of $5.0~\mathrm{BeV/c}$ K mesons at the Brookhaven AGS. About 100,000 pictures were taken, containing a total K track. The analysis led to a single event associated with the $\Omega^{-}\to nK^{0}\pi^{-}$ decay.. That extraordinarily good agreement provided strong support and confidence for the quark model of hadrons.
@@
Independent of Gell-Mann, Susumo Okubo (1930-2015) was pursuing his own studies of unitary symmetry in strong interactions at Rochester, leading to mass relations of the same type [48, 56, 57]. Since then, the SU(3) mass relations are known as Gell-Mann-Okubo (GMO) mass formulas [^9] Deviations from the GMO mass formula predictions often provide important hints for mass shifts caused by mixing effects from dynamical self-energy..
The interpretation, prediction, and description of hadron spectra in terms of constituent quarks paved the way for QCD. It was clear that Gell-Mann’s quark model could not be the final answer to strong interaction physics. A missing link was the lack of a theory incorporating the forces between quarks. In 1973, Harald Fritzsch, Murray Gell-Mann, and Heinrich Leutwyler published a letter [58] in which they presented the principles of a quantum field theory of strong interactions, pointing out the Advantages of the Color Octet Gluon Picture already in the title of their article.
A defining feature of QCD is the SU(3) color group, physically represented by three color charges. Color confinement had to be imposed in order to account for the empirical observation that hadrons exist only in color singlet states, i.e., they are colorless.
A central role is played by the spontaneously broken chiral symmetry, which leads to a complex, non-perturbative QCD vacuum composed of gluon and quark condensates. The pseudo-scalar mesons shown below in Fig. 7 represent, for example, the Goldstone bosons of broken chiral symmetry in the $u,d,s$ sector.
Figure 5: Properties of the current quarks as appearing in the QCD Lagrangian of the first $(u,d)$, second $(c,s)$, and third $(t,b)$ generation [41]. As a generalization of isospin, originally introduced for $\{u,d\}$ systems and used above to distinguish proton and neutrons by the eigenvalues of the third component of the isospin operator $I_{3}=\pm\frac{1}{2}$, heavier quarks are characterized by the flavors $c,s,t,b$ with values $0,1$. In this chapter, only production processes involving hadrons built of the first and second quark generations are considered. The figure is adapted from Ref. [59].
The works of Yuval Ne’eman, Murray Gell-Mann, and George Zweig were the long-awaited breakthroughs towards a new understanding of hadrons in terms of a few elementary degrees of freedom. The theory of Fritzsch, Gell-Mann, and Leutwyler provided the missing links for a full-scale theory of strong interaction physics. The modern understanding of hadrons and their interactions, emerging from those ground-breaking works, is centered around elementary massive quark fermion fields, interacting by gluon gauge vector fields as the force carriers of strong interactions. Different from the photons of QED and the weak interaction gauge bosons $Z^{0},W^{\pm}$, gluons carry color charges and interact with each other. QCD is a non-Abelian quantum gauge field theory of highly non-linear character. The complexities manifest in a rich spectrum of particles, but also in peculiarities such as running coupling constants, color confinement, and quarks with fractional baryon and charge numbers.
Long before the advent of the SM, Peter Higgs (1929-2024) proposed in 1964 a general mechanism for mass generation, including a multiplet of bosons [^10] Particles with integer spin are named bosons in honor of Satyendra Nath Bose (1894-1974).. At the same time, François Englert (born 1932) and Robert Brout (1928-2011) independently discovered the same mechanism. After an initial period of ignorance, in the late 1960s, the community became aware of the deep relevance of that mechanism and adopted it informally as Higgs mechanism. In 2012, indeed, a boson of about the right mass was found at CERN [60, 61]. Further studies confirmed that this was, in fact, the long-sought (neutral) Higgs boson $H^{0}$ (colloquially called the “God Particle”) of mass $m_{H}=(125.20\pm 0.11)~\mathrm{GeV}$ [41]. Higgs and Englert received the Nobel Prize in 2013 for the theoretical discovery of a mechanism that contributes to our understanding of the origin of the mass of subatomic particles, which was recently confirmed through the discovery of the predicted fundamental particle by the ATLAS and CMS experiments at CERN’s Large Hadron Collider.
With the discovery of the Higgs boson, the Standard Model (SM) of elementary particle physics has achieved the status of an intrinsically closed theory. The combined predictive power of QCD and electroweak theory is confirmed in all experimental tests. Moreover, until now, the SM has survived all attempts to search for physics beyond the SM.
Electroweak theory was developed in the 1960s by Sheldon Glashow (born 1932), Mohammad Abdus Salam (1926-1996), and Steven Weinberg (1933-2021). They received equal parts of the 1979 Nobel Prize in Physics for their contributions to the theory of the unified weak and electromagnetic interactions between elementary particles, including, inter alia, the prediction of the weak neutral current.
2 From QCD to Hadrons
Hadron physics investigates condensed asymptotic mass eigenstates of QCD, held together by the strong forces among quarks and gluons, and additionally stabilized by the surrounding polarization cloud. Hadrons are a generic new form of bound states. Attempts to separate the constituents fail, except under the extreme conditions of a highly compressed and heated quark-gluon plasma. Spectroscopic studies are demanding due to their reliance on powerful accelerators and detector systems.
However, attempts to explore the interior of hadrons have been made since the beginning of modern elementary particle physics. First access to the intrinsic properties of the nucleon was gained as early as 1955 by Robert Hofstadter (1915-1990). Hofstadter was a pioneer in electron-nucleus scattering and used the same methods to measure the electromagnetic form factors of the proton [62, 63]. Those measurements and the neutron data of Sobottka et al. [64] finally led to two revolutionary conclusions: Firstly, the nucleon is an extended object. Secondly, protons and neutrons both have a charged core, which, in a proton, is surrounded by a cloud of neutral pions, while in neutrons, the internal charge is compensated by a cloud of oppositely charged pions. Hofstadter shared with Mössbauer the Nobel Prize in 1961 for his pioneering studies of electron scattering in atomic nuclei and for his subsequent discoveries concerning the structure of nucleons.
On the energy scale of modern hadron physics, the Hofstadter experiments, using electron beams of about $100~\mathrm{MeV}$, were conducted at the lowest limit of quasi-elastic scattering with limited resolving power, where the collective features of the target prevail. About a decade later, when beams of much higher energy were available, deep-inelastic scattering revealed many more details of nucleon structure, as was exploited by Friedman et al. around 1970 [65]. Friedman started his research in Hofstadter’s group at Stanford. He then began a collaboration with Kendall on probing nuclear targets with the $22~\mathrm{GeV}$ electron beam at SLAC in the late 1960s. The much increased resolving power is compatible with an internal three-particle structure of the proton, in perfect agreement with the constituent quark model. Jerome Isaac Friedman (born 1930) received the 1990 Nobel Prize in Physics, together with Henry Kendall (1926-1999) and Richard Taylor (1929-2018) for their pioneering investigations concerning the deep inelastic scattering of electrons on protons and bound neutrons, which have been of essential importance for the development of the quark model in particle physics.
@@ 2 From QCD to Hadrons
These works were key contributions to the still ongoing efforts to unravel the secrets of hadron structure. Hadrons are understood as confined, color-neutral systems of quarks, fermions of spin $s=\frac{1}{2}$ with fractional electric charge, flavor, and color, interacting via gluons, colored spin-$1$ and self-interacting vector gauge fields. Renormalization group techniques and regularization methods are indispensable components of the QCD mathematical toolbox. Spontaneous chiral symmetry breaking is a defining mechanism of QCD. The complexities of the non-Abelian structures of QCD lead to higher-order non-linear self-interactions, which inhibit solving the deeply nested field equations in closed form. Only in the ultra-high energy regime of asymptotic freedom beyond about $250~\mathrm{GeV}$, where perturbative treatments become possible and leading-order results can be derived analytically, was the work of David J. Gross (born 1941), H. David Politzer (born 1949) and Frank Wilczek (born 1951) honored with the Nobel Prize (in 2004) for their discovery of asymptotic freedom in the theory of the strong interaction.
@@
Figure 6: The first two baryon SU(3) flavor multiplets are given by an octet (left) and a decuplet (right). The valence quark content of the baryons is indicated explicitly. The vertical axes represent the hypercharge $Y=S+B$, given by the strangeness $S$ and baryon number $B$; the horizontal axes indicate the third component $I_{3}=Q-\frac{1}{2}Y$ of the isospin $I$, which also includes the charge number $Q$. The group theoretical background and construction of these kinds of diagrams are discussed in depth in textbooks, see e.g.[36].
@@
Mesons and baryons are the asymptotic mass eigenstates of QCD, but they are fundamentally different in character. Considered as systems of valence quarks and gluons, mesons are formed by an even number of valence quarks, while baryons are defined as particles containing an odd number of valence quarks. In both cases, an arbitrary number of gluons is allowed in addition. Since quarks are Fermions with spin $s=\frac{1}{2}$ and positive parity, baryons carry half-integer spin quantum numbers $J_{B}=\frac{1}{2},\frac{3}{2},\frac{5}{2},\ldots$, while mesons have integer spin quantum numbers $J_{M}=0,1,2\ldots$. Moreover, baryons play a special role in cosmology because they account for protons and neutrons, along with a minor contribution from electrons, in the matter content of the Universe.
@@
Figure 7: The lowest meson SU(3) flavor nonets of pseudo-scalar (left), vector (center), and scalar mesons (right) as functions of hypercharge Y and isospin I3. Different to the baryons, mesonic two-quark systems of $\{u,d,s\}\otimes\{\bar{u},\bar{d},\bar{s}\}$ structure arrange into SU(3) octets and accompanying singlets. The isoscalar (charge-neutral) octet mesons obtain an isoscalar singlet partner. The singlet mesons are included into the octet diagrams, giving rise to the octet-singlet pairs $\{\eta,\eta^{\prime}\}$, $\{\omega,\phi\}$, and $\{\sigma\simeq f_{0}(500),\sigma^{\prime}\simeq f_{0}(980)\}$ in the pseudo-scalar, vector, and scalar meson sectors, respectively. Singlets and octets are mixed by, in principle, arbitrary mixing angles. A widely used choice is Ideal Mixing, chosen such that the isoscalar octet mesons $\eta,\omega,\sigma$ do not contain a $s\bar{s}$ component. Electromagnetic effects lead to $\omega-\rho$ mixing, and correspondingly, $\sigma-\delta^{0}$ mixing has to be expected. Scalar mesons may contain multi-pion, i.e., multi-$q\bar{q}$ components. The complexities of scalar mesons are especially visible in the isovector scalar meson $\delta\simeq a_{0}(980)$. $a_{0}(980)$ is decaying preferentially into $\eta\pi$, finally ending in three pions, but also $K\bar{K}$ and $\eta^{\prime}\pi$ decays are reported [41]. The physical counterpart of the $\kappa$ meson is the $K^{\ast}_{0}(700)$ meson of a notoriously uncertain pole structure [41]. Thus, the physically observed mesons are not identical with the bare SU(3) mesons.